Language: ngerman

today  normal date format 16. Juni 2000
isodate  2000-06-16
numdate  16. 06. 2000
shortdate  16. 06. 00
TeXdate  2000/06/16
back to normal  16. Juni 2000
another day normal date format 30. Januar 1972
isodate  1972-01-30
numdate  30. 01. 1972
shortdate  30. 01. 72
TeXdate  1972/01/30
back to normal  30. Januar 1972

Language: austrian

today  normal date format 16. Juni 2000
isodate  2000-06-16
numdate  16. 06. 2000
shortdate  16. 06. 00
TeXdate  2000/06/16
back to normal  16. Juni 2000
another day normal date format 30. Jänner 1972
isodate  1972-01-30
numdate  30. 01. 1972
shortdate  30. 01. 72
TeXdate  1972/01/30
back to normal  30. Jänner 1972

Language: english

today  normal date format 16th June 2000
isodate  2000-06-16
numdate  16/06/2000
shortdate  16/06/00
TeXdate  2000/06/16
back to normal  16th June 2000
another day normal date format 30th January 1972
isodate  1972-01-30
numdate  30/01/1972
shortdate  30/01/72
TeXdate  1972/01/30
back to normal  30th January 1972

Language: american

today  normal date format June 16, 2000
isodate  2000-06-16
numdate  06/16/2000
shortdate  06/16/00
TeXdate  2000/06/16
back to normal  June 16, 2000
another day normal date format January 30, 1972
isodate  1972-01-30
numdate  01/30/1972
shortdate  01/30/72
TeXdate  1972/01/30
back to normal  January 30, 1972

Changed ISO output

2000-12-24 2000–12–24 2000·12·24